# Learning 3D Object Spatial Relationships from Pre-trained 2D Diffusion Models

### v1

@@ 3.3 OOR Diffusion
Text Context Augmentation. For better generalization, we leverage an LLM [42] to diversify the OOR context $\mathbf{c}$ in two ways. First, we augment $\mathbf{c}$ with variations in prompt phrasing (e.g., verbs and sentence structure), while preserving the original semantic meaning. Second, we also diversify the object categories where the same relations hold. For example, the context “A teapot pours tea into a teacup” can represent similar spatial relations with “A kettle pours tea into a cup”, despite referring to different object pairs. We prompt the LLM to suggest alternative object categories with similar shapes and sizes that can substitute the context $\mathbf{c}$, while preserving the same OOR distributions. We find that our text context augmentation improves generalization by enabling OOR modeling under a broader range of contextual conditions. For details on data augmentation, see Sec. B. As a result, we train our model for 475 distinct contexts with 23750 text prompts, each capturing various spatial relationships between object pairs across 188 object categories.
@@
To address these issues, we present an approach to infer all multi-object OORs $\Phi=\left\{\phi_{t}^{p_{i}}\right\}_{i=1}^{n}$ simultaneously, by including our novel loss terms: collision loss $C(\Phi)$ and inconsistency loss $I(\Phi)$. This modifies the reverse ODE process in Eq. 9 for all $i$ as follows:
@@ 4 Experiments
We demonstrate the practical efficacy and generalization of our method through four main experiments. In Sec. 4.1, we show that our diffusion model generates OORs that fit the text context more plausibly and effectively than other methods. Sec. 4.2 demonstrates our advanced sampling approach produces significantly better results compared to text-to-3D models. In Sec. 4.3, we present an optimization technique that leverages the characteristics of the score-based model as an application of our OOR diffusion. Finally, Sec 4.4 presents an ablation study, verifying the superiority of our dataset generation method.
@@ 4.1 Pairwise OOR Generation
Baseline Methods. Since the concept of object-object spatial relationships (OOR) is new in the field, there exists no existing competitor with the same goal. Thus, we compare our method with LLM-based methods that achieve similar goals: SceneMotifCoder(SMC) [61], which focuses on 3D object alignments, and SceneTeller [41], which deals with 3D layout on a plane.
SceneMotifCoder (SMC) [61] is an example-driven visual program learning method. It takes text prompts as input and produces 3D object alignments by selecting and arranging meshes from a mesh pool. Given an example of a GT text prompt and mesh alignments in training, SMC analyzes patterns within the input, generates programs, and updates the program when new examples are introduced. During the inference process, when an input text comes in, LLM maps it to an appropriate task, and the program produces 3D object arrangements with the retrieved meshes from the candidate mesh pool. We convert our template mesh pairs obtained in the dataset generation process (Sec. 3.2) into the SMC format. Since SMC is not concerned about the relative scale between objects, we use meshes with scales in our OOR dataset as a mesh pool during inference.
@@
Table 1: Quantitative Comparisons of Pairwise OOR Generation. For each method, we evaluate the CLIP score [46], the VQA score [34], our proposed VLM score [68], and conduct a user study.
Metrics. Since there is no existing metric for OOR, we introduce several evaluation metrics to assess text prompt alignment with the plausibility of spatial relationship outputs in multi-view renderings. The CLIP score [46] measures text-image alignment by averaging CLIP model logits. The VQA score [34] leverages a VQA model to assess object composition and relations. However, since these methods evaluate per single image, we adopt GPTEval3D [68] to evaluate the text-to-3D alignment in multi-view images by leveraging VLMs [42]. Inspired by this, we propose the VLM score, which evaluates how well OOR is represented in multi-view images. The VLM score is measured in a pairwise manner, averaging the percentage of times a method is preferred (with a maximum score of 100), while allowing ties. Additionally, we conduct a user study on CloudResearch, surveying a total of 92 respondents. Details on guided prompts for VLM and the questionnaire for user study are in Sec.  C.
Results. We evaluate both our method and the baselines on a total of 150 scenes derived from 30 category pairs with 5 scenes generated per prompt. Each scene is rendered into $10$ views with a white background. SMC provides the arranged meshes itself, while SceneTeller provides the relative pose and scale. The baselines can reason about simple spatial relationships such as “on” or “next to” reasonably but struggle with functional relationships. Fig. 6 shows a qualitative comparison for rendered multi-view images between our method and baselines for these examples. In the case of SMC, despite the program handling translation appropriately, it often completely misinterprets rotation information. SceneTeller, on the other hand, benefits from the strong in-context learning capability of LLMs, enabling reasonable estimations. However, due to the inherent limitation of estimating 3D information without direct 3D data, it lacks fine-grained control. In contrast, our OOR diffusion demonstrates superior sampling capabilities compared to the baselines, leveraging its effective learning of the OOR distribution with respect to the text context. Tab. 1 shows that our method outperforms baselines for all metrics.
@@
Table 2: Quantitative Comparisons of Multi-object OOR Generation.
4.2 Multi-object OOR Generation
Baseline Method. Since SMC and SceneTeller cannot be directly extended to multi-object OOR using only pairwise OOR data, we compare our model to another baseline GraphDreamer [13], which generates compositional 3D scenes from text inputs. While GraphDreamer generates multiple objects, it employs neural implicit representation [66] for 3D objects, which may result in low texture quality. This difference makes direct comparison challenging, as our method renders scenes with template meshes. To mitigate the impact of rendering quality, we use ControlNet [74] to standardize the output style of synthesized images by using Canny edge or depth as guidance.
@@ 4.2 Multi-object OOR Generation
4.3 Applications of OOR for Scene Editing
Our OOR diffusion, built on a score-based model, enables the application of optimization techniques in various 3D scene editing scenarios.
Denoising 3D Scene Arrangement. Given a scene with multiple objects in a noisy arrangement, we use our method to “denoise” the arrangement, producing a more plausible and realistic object arrangement. Intuitively, we use our OOR diffusion to guide the noisy input arrangements toward high-likelihood regions in the OOR distribution. For simplicity, consider a scene with a pair of objects, a base object and a target object, $(\mathcal{B},\mathcal{T})$, with a relationship defined by context $\mathbf{c}$. Given a noisy scene arrangement $\phi_{0}$ as input, our method optimizes the OOR $\phi\sim p_{\mathbf{c}}^{\mathcal{T}\rightarrow\mathcal{B}}$ by minimizing the following objective function:
$\begin{gathered}\mathcal{L}_{\text{arrange}}(\phi)=||\phi-\phi_{0}||^{2}_{2}-\lambda_{1}\log p_{\mathbf{c}}^{\mathcal{T}\rightarrow\mathcal{B}}(\phi)\end{gathered},$ (11)
@@ 4.3 Applications of OOR for Scene Editing
where $\eta=0.01$ and $\lambda_{1}=0.01$. Since the scale of each object is predetermined here, we only need to update the relative pose $\phi^{\prime}=(\mathbf{R}^{\mathcal{T}\rightarrow\mathcal{B}},\mathbf{t}^{\mathcal{T}\rightarrow\mathcal{B}})$. For multi-object scenarios, we leverage the introduced loss terms, $C(\Phi)$ and $I(\Phi)$, in Eq. 10. In practice, optimization is completed within $50$ steps.
Applying Different Contexts for the Rearrangement. We use our method to apply different scene contexts $\mathbf{c}$ given a scene arrangement. This can be performed by replacing the input context $\mathbf{c}$ in our OOR diffusion and applying the rearrangement with our optimization pipeline (Eqs. 11– 13).
Adding New Objects. Our method enables adding new objects into the scene by applying multi-object OOR generation (Eq. 10) to optimize their relative poses, while keeping the OORs in the existing scene and the scale of the inserted objects fixed.
@@
Results. We demonstrate the efficacy of our OOR models for scene arrangement tasks using ParaHome DB [27], which provides 3D scenes with separate object meshes. Fig. 8(a) shows the result of rearranging (denoising) a scene after adding random noise to the original scene. In Fig. 8(b), we apply a scene context, where the teapot, which is initially placed next to the teacup, is adjusted to pour tea into the teacup. Finally, in Fig. 8(c), we demonstrate adding a pan and a salt shaker which are not originally present in the table, and then applying an optimization with a different scene context: “A salt shaker sprinkles salt into a pan”.
@@
4.4 Ablation Study
Baseline Methods. We ablate our mesh registration pipeline (Sec. 3.2) by removing point cloud separation and PCA on semantic features (SD+DINO [72]), comparing them with our full pipeline.
@@ 4.4 Ablation Study
5 Discussion
We present a novel approach to learning object-object spatial relationships (OOR) from synthetic 3D samples, leveraging pre-trained 2D diffusion models. We formulate a novel concept of OOR, and introduce a comprehensive pipeline for synthesizing 3D OOR samples, which can be applicable for any unbounded object categories. We develop a score-based diffusion model specifically designed to model 3D spatial relationships between objects. Furthermore, we demonstrate the scalability of our method by proposing a method to generate muti-object OOR by extending our pairwise OOR models. We also demonstrate the potential of our method for real-world tasks by showing that scene editing can be effectively implemented via score-based updates. As a future direction, more factors, such as detailed object shapes, can be considered as additional factors to determine OOR.
Acknowledgements. We thank Inhee Lee and Taeksoo Kim for their assistance in conducting the user study. This work was supported by RLWRLD, NRF grant funded by the Korean government (MSIT) (No. 2022R1A2C2092724), and IITP grant funded by the Korea government (MSIT) [No. RS-2024-00439854, RS-2021-II211343, and No.2022-0-00156]. H. Joo is the corresponding author.
Appendix A Details on Dataset Generation
Filtering Process. Our filtering process is automated in the following steps: First, we filter out all the bad quality images where segmentation and SfM fail. In the process from SfM to feature matching, we filter out if the number of points corresponding to each base object and target object is less than $100$. The cosine similarity threshold is set to $0.7$ in most cases. There may still be misalignment between the registered mesh and the point clouds. We use the Chamfer Distance from the mesh to the point clouds. The threshold is adjusted according to the scale of the registered mesh. Most of the bad samples are filtered out through a series of processes, but some cases, such as flipped meshes, may remain. We use VLM [42] to filter out the last remaining bad samples. Specifically, we render combinations of base and target objects, and then ask VLM to judge whether the multi-view images align well with the text prompt, using the same criteria as when measuring the VLM score. The filtering ratio is $0.46$ to $0.88$. However, since our approach is based on fully synthetic data, we can iterate this process as needed to obtain a sufficient number of high-quality 3D outputs. We obtain $30$ to $216$ samples per context.
Appendix B Details on OOR Diffusion.
Text Context Augmentation. As proposed in Sec. 3.3, we perform text context augmentation to increase the generality of OOR diffusion. Through the guided prompts in Fig. 12, LLM generates various text prompts that describe a given context $\mathbf{c}$. Object categories are augmented by asking the LLM to present categories with similar shape and scale that could replace $\mathcal{B}$ and $\mathcal{T}$ in the given text context $\mathbf{c}$.
Inconsistency Loss. The inconsistency loss introduced in Eq. 10 is computed as the average of the following three inconsistency parts: (1) The scale variance of a global base object; (2) The pose and scale variance in the global coordinate system derived from different parents; (3) The variance of each component’s ratio between the scale in the global coordinate system and the base scale in pairwise OOR, measured for parent nodes that are not global bases. Specifically, (1) corresponds to the part related to the desk in Fig. 5. OOR diffusion generates different $\mathbf{s}^{\mathcal{B}}$ for each pair, (desk, monitor), (desk, keyboard), and (desk, mouse), within the batch. In this case, (1) takes the variance of three $\mathbf{s}^{\mathcal{B}}$ as a loss term. (2) is the part corresponding to the keyboard in Fig. 5. The pose and scale of each object in the global coordinate system are obtained as many times as the number of parent nodes of the corresponding node in the scene graph. Thus, the variance of three different poses and scales in global coordinate systems of the keyboard obtained from the desk, monitor, and mouse is the loss term in (2). (3) relates to the scale ratio consistency of monitor and mouse, which are parent nodes but not the global base. For example, the monitor should maintain consistency between its scale in the global coordinate system obtained from paths in the scene graph and the $\mathbf{s}^{\mathcal{B}}$ of the (monitor, keyboard) OOR sample. They do not have to be equal, but the ratio of each component should be constant. For example, if the obtained monitor scale in the global coordinate system is $(0.5,0.4,0.2)$, then the $\mathbf{s}^{\mathcal{B}}$ in the (monitor, keyboard) OOR sample should be $(1.0,0.8,0.4)$. The variance of the ratio of each component of the relevant scales is the loss term in (3).
Appendix C Experiments Details
VLM Score. As described in Sec. 4.1, we propose the VLM score, inspired by GPTEval3D [68], to evaluate the alignment between the text context of OOR and multi-view images. We use VLM (specifically GPT-4o [42]) to compare two sets of multi-view images, each containing 10 images. These image sets are generated using our method or baselines. VLM is tasked with selecting the image set that better represents the spatial relationship between objects described in the text prompt. To ensure a fair comparison, we instruct VLM to ignore texture quality and focus solely on OOR. Fig. 14 illustrates the guided prompt we provided to VLM along with an example response from VLM.
@@ Appendix C Experiments Details
Appendix D Generality of Our methods
Fig. 15 demonstrates the generality of our OOR modeling. Our OOR diffusion generates appropriate relative poses and scales even for instances other than the template meshes used to generate the dataset. We consider the following scales to maintain the aspect ratio of each instance for both the base object and the target object:

### v2

@@ 3.3 OOR Diffusion
Text Context Augmentation. For better generalization, we leverage an LLM [41] to diversify the OOR context $\mathbf{c}$ in two ways. First, we augment $\mathbf{c}$ with variations in prompt phrasing (e.g., verbs and sentence structure), while preserving the original semantic meaning. Second, we also diversify the object categories where the same relations hold and have the similar shape. For example, the context “A teapot pours tea into a teacup” can represent similar spatial relations with “A kettle pours tea into a cup”. We ask LLM to suggest alternative object categories with similar shapes and sizes, while preserving the same OOR distributions. Our text context augmentation improves generalization by enabling OOR modeling under a broader range of contextual conditions. For details on data augmentation, see Sec. B. As a result, we train our model for 475 distinct contexts with 23750 text prompts, each capturing various spatial relationships between object pairs across 188 object categories.
@@
To address these issues, we present an approach to infer all multi-object OORs $\Phi=\left\{\phi_{t}^{p_{i}}\right\}_{i=1}^{n}$ simultaneously, by including our novel inference-loss terms: collision loss $C(\Phi)$ and inconsistency loss $I(\Phi)$. This modifies the reverse ODE process in Eq. 9 for all $i$ as follows:
@@ 4 Experiments
In this section, we demonstrate the practical efficacy and generalization of our method. In Sec. 4.1, we show that our diffusion model generates OORs that fit the text context more plausibly and effectively than other methods. Sec. 4.2 demonstrates our advanced sampling approach produces significantly better results compared to text-to-3D models. In Sec. 4.3, we present several applications that utilize our OOR. These are based on an optimization technique that leverages the characteristics of the score-based model.
@@ 4.1 Pairwise OOR Generation
Baseline Methods. Since the concept of OOR is new in the field, there is no existing competitor with the same goal. Thus, we compare ours with LLM-based methods that pursue similar goals: SceneMotifCoder (SMC) [60], which focuses on 3D object alignments, and SceneTeller [40], which deals with 3D layout on a plane.
@@
Table 1: Quantitative Comparisons of Pairwise OOR Generation. For each method, we evaluate CLIP score [46], VQA score [34], our proposed VLM score [67], and user study.
Metrics. Since there is no existing metric for OOR, we introduce several evaluation metrics to assess text prompt alignment with the plausibility of spatial relationship outputs in multi-view renderings. The CLIP score [46] measures text-image alignment by averaging CLIP model logits. The VQA score [34] leverages a VQA model to assess object composition and relations. However, since these methods evaluate per single image, we adopt GPTEval3D [67] to evaluate the text-to-3D alignment in multi-view images by leveraging VLMs [41]. Inspired by this, we propose the VLM score, which evaluates how well OOR is represented in multi-view images. The VLM score is measured in a pairwise manner, averaging the percentage of times a method is preferred (with a maximum score of 100), while allowing ties. Additionally, we conduct a user study on CloudResearch, surveying a total of 92 respondents.
Results. We evaluate our method and the baselines on a total of 150 scenes derived from 30 category pairs with 5 scenes generated per prompt. Each scene is rendered into $10$ views with a white background. SMC outputs the arranged meshes, while SceneTeller provides the relative pose and scale. Both baselines handle simple relations such as “on” or “next to” reasonably, but they struggle with functional relationships. As shown in Fig. 6, SMC often completely misinterprets rotation information despite handling translation appropriately. On the other hand, SceneTeller benefits from the strong in-context learning capability of LLMs, enabling reasonable estimations. However, due to the inherent limitation of estimating 3D information without direct 3D data, it lacks fine-grained control. In contrast, our OOR diffusion demonstrates superior sampling capabilities compared to the baselines, leveraging its effective learning of the OOR distribution with respect to the text context. Tab. 1 shows that our method outperforms baselines for all metrics.
@@
4.2 Multi-object OOR Generation
Baseline Method. Since SMC and SceneTeller cannot be directly extended to multi-object OOR using only pairwise OOR data, we compare our model to another baseline GraphDreamer [14], which is diffusion-based text-to-3D model. While GraphDreamer employs neural implicit representation [65] for 3D objects, which may result in low texture quality. Since our method renders scenes with template meshes, direct comparison is challenging. To mitigate the impact of rendering quality, we use ControlNet [73] to standardize the output style of synthesized images by using Canny edge or depth as guidance.
@@ 4.2 Multi-object OOR Generation
Table 2: Quantitative Comparisons of Multi-OOR Generation.
@@
4.3 Applications of OOR
3D Scene Editing. Our OOR diffusion, built on a score-based model, enables the application of optimization techniques in various 3D scene editing scenarios. First, we present a method that denoises a scene with multiple objects in a noisy arrangement, producing a more plausible and realistic configuration by guiding it toward high-likelihood regions in the OOR distribution. For simplicity, consider a scene with a pair of objects, a base object and a target object, $(\mathcal{B},\mathcal{T})$, with a relationship defined by context $\mathbf{c}$. Given a noisy scene arrangement $\phi_{0}$ as input, our method optimizes the OOR $\phi\sim p_{\mathbf{c}}^{\mathcal{T}\rightarrow\mathcal{B}}$ by minimizing the following objective function:
[…]
@@ 4.3 Applications of OOR
where $\eta=0.01$ and $\lambda_{1}=0.01$. Since the scale of each object is predetermined here, we only need to update the relative pose $\phi^{\prime}=(\mathbf{R}^{\mathcal{T}\rightarrow\mathcal{B}},\mathbf{t}^{\mathcal{T}\rightarrow\mathcal{B}})$. For multi-object scenarios, we leverage the introduced loss terms, $C(\Phi)$ and $I(\Phi)$. In practice, optimization is completed within $50$ steps. Also, we use our proposed optimization to apply different scene contexts $\mathbf{c}$ to a given arrangement. This can be performed by replacing the input context $\mathbf{c}$ in our OOR diffusion and optimizing with Eqs. 11– 13. Moreover, we add new objects into the scene by performing multi-object OOR generation (Eq. 10) while keeping the OORs in the existing scene and the scale of the inserted objects fixed.
@@
We demonstrate the efficacy of our proposed scene arrangement methods using ParaHome DB [28], which provides 3D scenes with separate object meshes. Fig. 8(a) shows the result of rearranging (denoising) a scene after adding random noise to the original scene. In Fig. 8(b), we apply a scene context, where the teapot, which is initially placed next to the teacup, is adjusted to pour tea into the teacup. Finally, in Fig. 8(c), we demonstrate adding a pan and a salt shaker which are not originally present in the table, and then applying an optimization with a different scene context: “A salt shaker sprinkles salt into a pan”.
Figure 9: Human Motion Synthesis Results. We assume two objects and a human interacting initially, and synthesize human motion using the generated OOR and contact consistency.
Human Motion Synthesis for Two-Object Interaction. As mentioned in Sec. 1, OOR often reflects how humans interact with objects. We demonstrate this by synthesizing human motion interacting with two objects using generated OORs. First, we begin with the initial scene with two objects and the human interacting with them. For instance, a teapot and a teacup are placed on a desk, and the human has just started to grasp the teapot’s handle. By applying Eq. 13, we obtain the sequence of OOR $(\mathbf{R}^{\mathcal{T}\rightarrow\mathcal{B}},\mathbf{t}^{\mathcal{T}\rightarrow\mathcal{B}})$ that takes it from the initial state to the OOR context $\mathbf{c}$. In the above example, setting $\mathbf{c}=$ “A teapot pours tea into a teacup”, $\mathcal{B}=$ “teacup”, and $\mathcal{T}=$ “teapot” in Eq. 13 yields a sequence where the teapot, initially placed beside the teacup, moves to pour tea into it. Then, we can synthesize human motion by applying constraints to maintain initial contact between the human and the objects across entire OOR sequence. For each state in the OOR sequence, the human pose is obtained by optimizing the following objective function:
[…]
where $\theta_{\text{VP}}$ is a human body pose embedding of VPoser [44], $\mathbf{A}$ is an angle prior of VPoser, $\mathcal{H}$ is a human derived from $\theta_{\text{VP}}$, $\mathbf{C}$ is a contact constraint between $(\mathcal{H},\mathcal{B},\mathcal{T})$, and $\lambda_{i}$ are weight terms. $\mathbf{C}$ maintains the distances of the $N$ nearest vertex pairs between the two object meshes and the human mesh at the initial state throughout the entire OOR sequence. Fig. 9 shows human motion interacting with two objects synthesized by this optimization.
5 Discussion
We present a novel approach to learning object-object spatial relationships (OOR) from synthetic 3D samples, leveraging pre-trained 2D diffusion models. We formulate a novel concept of OOR, and introduce a comprehensive pipeline for synthesizing 3D OOR samples, which can be applicable for any unbounded object categories. We develop a score-based diffusion model specifically designed to model 3D spatial relationships between objects. Furthermore, we demonstrate the scalability of our method by proposing a method to generate multi-object OOR by extending our pairwise OOR models. We also demonstrate the potential of our method for several applications by showing that scene editing and human motion synthesis can be effectively implemented via score-based updates. As a future direction, more factors, such as detailed object shapes, can be considered as additional factors to determine OOR.
Acknowledgements
We thank Inhee Lee and Taeksoo Kim for their assistance in conducting the user study. This work was supported by RLWRLD, NRF grant funded by the Korean government (MSIT) (No. RS-2022-NR070498 and RS-2025-25396144), and IITP grant funded by the Korea government (MSIT) (No. RS-2024-00439854, No.RS-2021-II211343, No.RS-2025-25441838, and No.2022-0-00156). H. Joo is the corresponding author.
Appendix A Details on Dataset Generation
Filtering Process. Our filtering process is automated in the following steps: First, we filter out all the bad quality images where segmentation and SfM fail. In the process from SfM to feature matching, we filter out if the number of points corresponding to each base object and target object is less than $100$. The cosine similarity threshold is set to $0.7$ in most cases. There may still be misalignment between the registered mesh and the point clouds. We use the Chamfer Distance from the mesh to the point clouds. The threshold is adjusted according to the scale of the registered mesh. Most of the bad samples are filtered out through a series of processes, but some cases, such as flipped meshes, may remain. We use VLM [41] to filter out the last remaining bad samples. Specifically, we render combinations of base and target objects, and then ask VLM to judge whether the multi-view images align well with the text prompt, using the same criteria as when measuring the VLM score. The filtering ratio is $0.58$ to $0.92$. However, since our approach is based on fully synthetic data, we can iterate this process as needed to obtain a sufficient number of high-quality 3D outputs. We obtain $30$ to $216$ samples per context.
Appendix B Details on OOR Diffusion.
Text Context Augmentation. As proposed in Sec. 3, we perform text context augmentation to increase the generality of OOR diffusion. Through the guided prompts in Fig. 12, LLM generates various text prompts that describe a given context $\mathbf{c}$. Object categories are augmented by asking the LLM to present categories with similar shape and scale that could replace $\mathcal{B}$ and $\mathcal{T}$ in the given text context $\mathbf{c}$.
Inconsistency Loss. The inconsistency loss introduced in Sec. 3.3 is computed as the average of the following three inconsistency parts: (1) The scale variance of a global base object; (2) The pose and scale variance in the global coordinate system derived from different parents; (3) The variance of each component’s ratio between the scale in the global coordinate system and the base scale in pairwise OOR, measured for parent nodes that are not global bases. Specifically, (1) corresponds to the part related to the desk in Fig. 5. OOR diffusion generates different $\mathbf{s}^{\mathcal{B}}$ for each pair, (desk, monitor), (desk, keyboard), and (desk, mouse), within the batch. In this case, (1) takes the variance of three $\mathbf{s}^{\mathcal{B}}$ as a loss term. (2) is the part corresponding to the keyboard in Fig. 5. The pose and scale of each object in the global coordinate system are obtained as many times as the number of parent nodes of the corresponding node in the scene graph. Thus, the variance of three different poses and scales in global coordinate systems of the keyboard obtained from the desk, monitor, and mouse is the loss term in (2). (3) relates to the scale ratio consistency of monitor and mouse, which are parent nodes but not the global base. For example, the monitor should maintain consistency between its scale in the global coordinate system obtained from paths in the scene graph and the $\mathbf{s}^{\mathcal{B}}$ of the (monitor, keyboard) OOR sample. They do not have to be equal, but the ratio of each component should be constant. For example, if the obtained monitor scale in the global coordinate system is $(0.5,0.4,0.2)$, then the $\mathbf{s}^{\mathcal{B}}$ in the (monitor, keyboard) OOR sample should be $(1.0,0.8,0.4)$. The variance of the ratio of each component of the relevant scales is the loss term in (3).
Appendix C Experiments Details
Baseline Methods Details for Pairwise OOR Generation. SceneMotifCoder (SMC) [60] is an example-driven visual program learning method. It takes text prompts as input and produces 3D object alignments by selecting and arranging meshes from a mesh pool. Given an example of a GT text prompt and mesh alignments in training, SMC analyzes patterns within the input, generates programs, and updates the program when new examples are introduced. During the inference process, when an input text comes in, LLM maps it to an appropriate task, and the program produces 3D object arrangements with the retrieved meshes from the candidate mesh pool. We convert our template mesh pairs obtained in the dataset generation process (Sec. 3.2) into the SMC format. Since SMC is not concerned about the relative scale between objects, we use meshes with scales in our OOR dataset as a mesh pool during inference.
VLM Score. As described in Sec. 4, we propose the VLM score, inspired by GPTEval3D [67], to evaluate the alignment between the text context of OOR and multi-view images. We use VLM (specifically GPT-4o [41]) to compare two sets of multi-view images, each containing 10 images. These image sets are generated using our method or baselines. VLM is tasked with selecting the image set that better represents the spatial relationship between objects described in the text prompt. To ensure a fair comparison, we instruct VLM to ignore texture quality and focus solely on OOR. Fig. 15 illustrates the guided prompt we provided to VLM along with an example response from VLM.
@@ Appendix C Experiments Details
Appendix D Ablation Study
Baseline Methods. We ablate our mesh registration pipeline in Sec. 3.2 by removing point cloud separation and PCA on semantic features (SD+DINO [71]), comparing them with our full pipeline.
Appendix E Generality of Our methods
Generality for Unseen Mesh Instances. Fig. 16 demonstrates the generality of our OOR modeling for unseen mesh instances. Our OOR diffusion generates appropriate relative poses and scales even for instances other than the template meshes used to generate the dataset. We consider the following scales to maintain the aspect ratio of each instance for both the base object and the target object: